OpenStudy (mwhit2101):

What's bigger? 4.72 or 4.6? and why?

OpenStudy (austinl):

Larger in the basic sense? \(\displaystyle 4.72=\frac{472}{100}\Rightarrow 4.6=\frac{460}{100}\) What would you say is bigger based on this? Or even by the decimals? It should be equally obvious.

OpenStudy (mwhit2101):

4.72 is larger?

OpenStudy (austinl):

Indeed.

OpenStudy (mwhit2101):

Can you help me with one more?

OpenStudy (austinl):

Sure :)

OpenStudy (mwhit2101):

is 0.5 larger or smaller than 0.5 (the second 0.5 has a bar over the 5 showing that it keeps going on)

OpenStudy (austinl):

\(0.5~vs.~0.\overline{5}\) \(0.\overline{5}\) means \(0.555555555555.....\) So, which one is larger?

OpenStudy (mwhit2101):

Well that's what I don't really know. It seems like 0.55555555555 should be bigger?

OpenStudy (dobby1):

it is

OpenStudy (dobby1):

.5555555555555555555555555555555555555555555555555555555555555555555555555 is much bigger

OpenStudy (austinl):

It would be! \(0.50~\text{versus}~0.55\) To clarify a little.

OpenStudy (mwhit2101):

oh ok. Thanks

OpenStudy (dobby1):

actuall .50 vs .56 if you simplify